OpenStudy (anonymous):

The Royal Fruit Company produces two types of fruit drinks. The first type is 45% pure fruit juice, and the second type is 95% pure fruit juice. The company is attempting to produce a fruit drink that contains 50% pure fruit juice. How many pints of each of the two existing types of drink must be used to make 120 pints of a mixture that is 50% pure fruit juice?

OpenStudy (kropot72):

Amount of pure fruit juice needed in 50% mixture = 120/2 = 60 pints Let x = amount of 45% juice needed. Then 120 - x = amount of 95% juice needed. Therefore the quantities of 45% and 95% juice drinks needed to make 60 pints of pure juice are expressed in the following equation: 0.45x + 0.95(120 - x) = 60 Now you just need to solve to find x. Can you do that?

OpenStudy (anonymous):

i already got it, but thanks anyway :)

OpenStudy (kropot72):

You're welcome.